Automobiles arrive at a toll plaza located at the entrance to a bridge at a rate of 51 per minute during the 5:00 to 6:00 p.m. hour (assume Exponential distribution).

  1. If an auto has just arrived, what is the probability that the next auto will arrive within 3 seconds?
    1. 1.00
    2. 0
    3. 0.284
    4. 0.922
Answer #1

Let X denotes the inter-arrival time between two arrivals of auto.

X ~ Exponential(51/60) or X ~ Exponential( 0.85)

The probability density function of X is

fx(z) = 0.85 * e-0.85.22 0

The cumulative density function of X is

Ex(z) = 1-e-0.85.2, > 0

If an auto has just arrived, what is the probability that the next auto will arrive within 3 seconds = 3/60 = 0.05 minutes

P(X < 0.05)

=1 - e^{-0.05*0.85}

0.0416095

ans-> B. 0